OpenStudy (anonymous):

what is the soultion of the following system {x-y=11 {-x+y=-11 A (-3,-4) B no solution C infinitely many solutions D (3,4)

OpenStudy (shamil98):

11 does not equal -11, so no solutions, the x and y values cancel each other out, which is how i arrived at this conclusion. No solution.

OpenStudy (anonymous):

thank you i thought that i was just making sure

ganeshie8 (ganeshie8):

wait a sec, it looks wrong x-y=11 -x+y=-11

OpenStudy (shamil98):

oh my bad

ganeshie8 (ganeshie8):

x-y=11 -x+y=-11 -------------- 0 = 0

OpenStudy (shamil98):

I mean infinitley many solutions ._. brain fart.

ganeshie8 (ganeshie8):

since left side and right side are equal, it will have infinite solutions !

OpenStudy (anonymous):

oh okay thank you

ganeshie8 (ganeshie8):

yw :)